The two crates m 2 = 40 kg and m 1 = 60 kg are to be accelerated upward to a ship at the rate of 0.30 m/s ^ 2 Determine the tension in each rope.
1
Expert's answer
2020-12-06T17:29:21-0500

According to the second Newton's law, the tension "T" in a cord is:


"T = ma"

where "m" is the mass of a crate and "a" is its acceleration. Thus, for the first crate have:


"T_1 =m_1a = 60kg\cdot 0.3m/s^2 = 18N"

And for the second


"T_2 =m_2a = 40kg\cdot 0.3m/s^2 = 12N"

Answer. "T_1 = 18N,\space T_2 = 12N."